OpenStudy (anonymous):

integrationn of cos2xsinx

OpenStudy (raden):

cosA sinB = 1/2(sin(A+B) - sin(A-B))

OpenStudy (anonymous):

One way would be ... write cos2x= 1-2 sin^2x ....

OpenStudy (raden):

int (cos2x sinx) dx = int (1/2(sin(2x+x) - sin(2x-x)) = 1/2 int (sin3x - sinx) dx = 1/2 int sin 3x dx - 1/2 int sinx dx does that helps ?

OpenStudy (shubhamsrg):

better thing might be to write cos2x = 2cos^2x - 1 and assume cos^2x =t

OpenStudy (shubhamsrg):

I mean assume cosx = t

OpenStudy (raden):

hint : use the basic formulas int sinx = -cosx and int sinAx = -1/A cosAx

OpenStudy (anonymous):

thank you :) it is helpful:D

OpenStudy (anonymous):

can you give me the final answer? i want to see if i correct:/

OpenStudy (raden):

well, actually that's just a basic integration problem :) 1/2 int sin 3x dx - 1/2 int sinx dx = 1/2(-1/3 cos3x) - 1/2(-cosx) + c = -1/6 cos3x + 1/2 cosx + c

OpenStudy (anonymous):

I just found the MS, the answer is \[-2/3\cos ^{3}x+cosx\]

OpenStudy (anonymous):

here is detailed solution

OpenStudy (anonymous):

thank you ~ maybe the MS is wrong:D